convert to rectangular form: r=3sin(theta)
Answer #1

I know x=r cos(theta).

Y=rsin(theta) .

r^2=x^2+y^2.

Now question is:

r=3sin(theta).

Multiply both sides by r.

r^2=9rsin(theta).

X^2+y^2=9y.

Hence rectangle form:

X^+y^2-9y=0